# assumption[theorem]Assumption

### v1

@@ Abstract
We prove a Tauberian transfer principle showing that for any compact closed Riemannian manifold $(M^{d},g)$ the affine spectral encoding $C=\pi-\epsilon\lambda$ transports Laplacian Weyl asymptotics to a Weyl law in the edge variable in the bulk regime $C\to$-$\infty$ (equivalently, $(\pi-C)/\epsilon\to\infty$): $N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}$ and $\rho_{\mathrm{bulk}}(C)\sim\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{(d-2)/2}$, so that $d$ and the Weyl constant $\gamma_{d}$ are recoverable from one–dimensional edge–variable data. Conversely, a bulk power law $N_{\mu_{C}}(C)\ \\ \sim\!A(\pi-C)^{\alpha}$ as $C\to$-$\infty$ implies $d=2\alpha$ and $\gamma_{d}=A\,\epsilon^{d/2}$. We establish the uniqueness of the affine rule among polynomial–type encodings $g(\lambda)=a-b\lambda^{k}L(\lambda)$ (the edge–variable exponent forces $k=1$) and stability under small perturbations $C=\pi-\epsilon\lambda+\delta(\lambda)$ with $\delta(\lambda)=o(\lambda)$. For constant–curvature model spaces we record strengthened correspondences for heat traces and spectral zeta, $H_{\mathrm{edge}}(s)=\Theta_{\Delta}(\epsilon s)$ and $\zeta_{\mathrm{edge}}(u)=\epsilon^{-u}\zeta_{\Delta}(u)$, and we realize multiplicities via generalized one-dimensional models (Krein strings). When a Weyl remainder $O(\Lambda^{(d-1)/2})$ is available, it transfers to a bulk remainder $O((\pi-C)^{(d-1)/2})$ in the CC–variable.
@@
weyl law; bulk-edge asymptotics; counting function; affine spectral encoding; laplace–beltrami operator; spectral zeta function; tauberian theorems; stability
@@ 1 Introduction
Theorem 1.1 (Main Theorem).
Let $(M^{d},g)$ be a smooth compact Riemannian manifold without boundary, and let $C=\pi-\epsilon\lambda$ be the affine encoding of the Laplace spectrum. Then:
(i) A Weyl law in the edge variable holds in the bulk regime $C\to$-$\infty$ (equivalently $(\pi-C)/\epsilon\to\infty$), yielding
@@
so that $d$ and $\gamma_{d}$ are recoverable from one–dimensional edge–variable data.
(ii) Among polynomial-type encodings $g(\lambda)\sim a-b\lambda^{k}L(\lambda)$ with $L\in\mathrm{RV}_{0}$, only the asymptotically affine case $k=1$ preserves the geometric edge–variable density exponent (d2)/2(d-2)/2.
(iii) The edge–variable exponents are stable under perturbations $C=\pi-\epsilon\lambda+\delta(\lambda)$ with $\delta(\lambda)=o(\lambda)$.
@@
The statement above rigorously realizes a principle informally articulated in our earlier work, where the edge–side spectral density was proposed as a dimensional signature (sec. VI.A in [1]). The present paper proves this mechanism via an affine pushforward of the Laplace spectrum and Tauberian tools, and further establishes the uniqueness of the affine encoding among polynomial–type rules and stability under $o(\lambda)$ perturbations.
@@
We reserve edge for the geometric endpoint $C\uparrow\pi$, while bulk refers to $C\to$-$\infty$. The asymptotics transferred from Weyl’s law occur in the bulk regime.
@@ 2.2 Spectral measures and counting
$N_{\Delta}(\Lambda):=\mu_{\Delta}((-\infty,\Lambda])=\#\{n:\lambda_{n}\leq\Lambda\},$ (2.3)
where multiplicities are counted [4, Thm. VII.2].
@@
Since $N_{\Delta}(\Lambda)=\mu_{\Delta}((-\infty,\Lambda])$ with a purely atomic measure, all stated properties are immediate.
@@ 2.3 Affine edge encoding and edge variable
Immediate from $dC/d\lambda=-\epsilon<0$.
@@ 2.4 Exact composition identity
Since $C=\pi-\epsilon\lambda$ is strictly decreasing, $\{n:C_{n}\geq C\}=\{n:\lambda_{n}\leq(\pi-C)/\epsilon\}$, with multiplicities.
@@ 2.6 Standing assumptions
with the classical remainder $O(\Lambda^{(d-1)/2})$ when available; [6, 7] and, for refinements and boundary effects, [8, Ch. 1];
@@
3 Edge-Variable Weyl transfer: main results
@@ 3.1 Edge-variable Weyl (Abelian direction)
By smoothing on scales $h=o(y)$ (Prop. 3.18) we may assume eventual monotonicity required by the monotone density theorem; the asymptotics are unchanged. By Karamata’s differentiation (monotone density) theorem for regularly varying functions (Bingham–Goldie–Teugels [9, §§1.4–1.7, Thm. 1.7.2]), if $F(y)=y^{\alpha}\ell(y)$ with $\alpha>0$ and $\ell\in\mathrm{RV}_{0}$ is ultimately monotone, then $F^{\prime}(y)\sim\alpha y^{\alpha-1}\ell(y)$. Hence
@@
so necessarily $\alpha=d/2$ and $\gamma_{d}=A\,\epsilon^{d/2}$.
@@
with ultimate monotonicity (or after smoothing as in Prop. 3.18), then
@@
Equivalently, the log–log slope $\alpha=\lim_{y\to\infty}\frac{d\log N_{\mu_{C}}(\pi-y)}{d\log y}$ yields $d=2\alpha$, and
@@
The same holds if one uses the (smoothed/Stieltjes) bulk density $\rho_{\mathrm{bulk}}(C)\sim B\,y^{\alpha-1}$ with $y=\pi-C$, where $d=2\alpha$ and $B=\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}$.
@@ 3.2 Weyl remainder in the edge variable (bulk regime)
If $N_{\Delta}(\Lambda)=\gamma_{d}\Lambda^{d/2}+O(\Lambda^{(d-1)/2})$ as $\Lambda\to\infty$ (see [6, 7, 8]), then, as $C\to$-$\infty$ (equivalently, $\Lambda=(\pi-C)/\epsilon\to\infty$),
@@
Apply Lemma 2.7 with $\Lambda=(\pi-C)/\epsilon$ and propagate the remainder. The substitution $\Lambda=(\pi-C)/\epsilon$ and monotonicity preserve both the order and form of the remainder.
@@
since $N_{\Delta}(Y)\sim\gamma_{d}Y^{d/2}$.
@@
If $N_{\Delta}(\Lambda)=\gamma_{d}\Lambda^{d/2}+O(\Lambda^{(d-1)/2})$ as $\Lambda\to\infty$, then after standard smoothing (as in Prop. 3.18), the bulk density satisfies
@@
Apply the monotone density theorem to the remainder estimate from Proposition 3.4, using the same smoothing procedure as in Proposition 3.18.
3.3 Tauberian reconstruction (inverse direction)
We adopt the edge–counting convention
$N_{\mu_{C}}(C)\;:=\;\#\{\ell:\;C_{\ell}\geq C\}\qquad(C<\pi),$ (3.19)
which counts the atoms of $\mu_{C}$ to the right of $C$ (towards the edge $C=\pi$).
Lemma 3.10 (Change–of–variables dictionary).
For every $C<\pi$ one has the identity
$N_{\mu_{C}}(C)\;=\;\#\{\ell:\;C_{\ell}\in[C,\pi]\}\;=\;\#\{\ell:\;\lambda_{\ell}\leq\Lambda\}\;=\;N_{\Delta}(\Lambda),\qquad\Lambda=\frac{\pi-C}{\epsilon}.$ (3.20)
Proof 3.11.
Since $C_{\ell}=\pi-\epsilon\lambda_{\ell}$, the condition $C_{\ell}\in[C,\pi]$ is equivalent to $\pi-\epsilon\lambda_{\ell}\geq C$, i.e. $\lambda_{\ell}\leq(\pi-C)/\epsilon=\Lambda$. Counting with multiplicities gives the claim.
The lemma shows that the “bulk” regime $C\to$-$\infty$ is exactly the high–energy regime $\Lambda\to+\infty$. Hence Weyl’s law and the corresponding asymptotics for $N_{\mu_{C}}$ are equivalent.
Theorem 3.12 (Tauberian reconstruction of $d$ and $\gamma_{d}$).
Let $A>0$ and $\alpha>0$. The following are equivalent:
$\textup{(i)}\quad N_{\mu_{C}}(C)\ \sim\ A\,\epsilon^{-\alpha}\,(\pi-C)^{\alpha}\qquad(C\to-\infty),$
$\textup{(ii)}\quad N_{\Delta}(\Lambda)\ \sim\ A\,\Lambda^{\alpha}\qquad(\Lambda\to+\infty).$
In particular, under Weyl’s law $N_{\Delta}(\Lambda)\sim\gamma_{d}\,\Lambda^{d/2}$ one has
$N_{\mu_{C}}(C)\ \sim\ \gamma_{d}\,\epsilon^{-d/2}\,(\pi-C)^{d/2}\qquad(C\to-\infty),$ (3.21)
and conversely, observing the power–law exponent $\alpha$ and the coefficient $A$ of the normalized form $N_{\mu_{C}}(C)\sim A\,\epsilon^{-\alpha}(\pi-C)^{\alpha}$ determines
$d=2\alpha,\qquad\gamma_{d}=A.$ (3.22)
Proof 3.13.
By Lemma 3.10, $N_{\mu_{C}}(C)=N_{\Delta}\big((\pi-C)/\epsilon\big)$. Set $y:=\pi-C$ and $\Lambda:=y/\epsilon$. Then
$N_{\mu_{C}}(\pi-y)=N_{\Delta}(y/\epsilon).$ (3.23)
Thus $N_{\mu_{C}}(\pi-y)\sim A\,\epsilon^{-\alpha}y^{\alpha}$ as $y\to+\infty$ is equivalent to $N_{\Delta}(\Lambda)\sim A\,\Lambda^{\alpha}$ as $\Lambda\to+\infty$. The specialization with $\alpha=d/2$ and $A=\gamma_{d}$ is immediate.
Remark 3.14 (Practical estimators from bulk data).
Given samples of $N_{\mu_{C}}(C)$ for large negative $C$, the following consistent estimators follow from Theorem 3.12:
$\widehat{d}(C)\;=\;2\,\frac{d\log N_{\mu_{C}}(C)}{d\log(\pi-C)},\qquad\widehat{\gamma}_{d}(C)\;=\;\epsilon^{\widehat{d}(C)/2}\,\frac{N_{\mu_{C}}(C)}{(\pi-C)^{\widehat{d}(C)/2}}.$ (3.24)
In discrete data, replace the derivative by a local log–log slope. By Lemma 7.14, clustering contributes only bounded microscopic oscillations and does not affect the leading bulk asymptotics. See also Theorem 3.12 and Proposition 3.18.
Remark 3.15 (Edge vs. bulk).
The limit $C\to$-$\infty$ corresponds to $\Lambda\to+\infty$ (Weyl/high–energy, “bulk”), while the limit $C\uparrow\pi$ probes only finitely many low modes and is not governed by Weyl. To avoid confusion with the geometric “edge” $C\uparrow\pi$, we refer to $\rho_{\mathrm{bulk}}$ in the $C\to$-$\infty$ regime.

### v2

# Anton Alexa

@@ Abstract
We prove that the geometric Weyl bulk-density exponent (d2)/2(d-2)/2 rigidifies spectral encodings $C=\pi-\phi(\lambda)$ in the O-regularly varying class: the bulk power law forces $\phi\in\mathrm{RV}_{1}$ (asymptotic linearity). For polynomial-type encodings $C=\pi-\epsilon\lambda^{k}L(\lambda)$ with $L\in\mathrm{RV}_{0}$, this yields the unique admissible exponent $k=1$. The affine encoding then gives $N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}$ as $C\to$-$\infty$, allowing recovery of $d$ and $\gamma_{d}$ from bulk encoded data. This transfer is stable under perturbations $\delta(\lambda)=o(\lambda)$, with explicit slowly varying error control. We further formalize asymptotic spectral equivalence classes: if $\phi\in\mathrm{RV}_{k}$, the induced map scales asymptotic spectral dimension as $d_{\mathrm{as}}\mapsto d_{\mathrm{as}}/k$; hence dimension preservation is equivalent to $\phi\in\mathrm{RV}_{1}$, with strict affine normalization at first order when $L(\lambda)\to 1$.
@@
Weyl law, spectral encoding rigidity, affine rigidity, O-regular variation, Tauberian theorems, bulk–edge asymptotics, asymptotic spectral dimension, spectral equivalence classes, Kreĭn strings, stability
@@ 1 Introduction
We study the following rigidity question: which spectral encodings of the Laplace spectrum preserve the geometric Weyl density exponent?
Let $(M^{d},g)$ be a smooth compact Riemannian manifold without boundary, with Laplace eigenvalues $0=\lambda_{0}\leq\lambda_{1}\leq\cdots\to\infty$. A spectral encoding is a reparametrization $C=f(\lambda)$ that maps the spectrum to a new variable; the induced counting function $N_{\mu_{C}}$ and bulk density $\rho_{\mathrm{bulk}}$ carry whatever spectral information survives the reparametrization. For this framework to be useful in geometric inverse problems, the encoding must be rigid: it must neither destroy nor artificially create the exponent (d2)/2(d-2)/2 that signatures dimension in Weyl’s law [6, 7, 8].
Our main result establishes rigidity within the full class of O-regularly varying encodings [9] (Theorem 4.12): the bulk power-law constraint forces $\phi\in\mathrm{RV}_{1}$, i.e., asymptotic linearity of the encoding. In the polynomial case $C=\pi-\epsilon\lambda^{k}L(\lambda)$ with $L\in\mathrm{RV}_{0}$, this specializes to $k=1$: any $k\neq 1$ distorts the exponent to (d2)/(2k)(d-2)/(2k), breaking spectral-dimension recovery.
As a structural consequence of the uniquely correct affine encoding $C=\pi-\epsilon\lambda$, the edge-variable Weyl law
$N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2},\qquad C\to-\infty,$
holds, allowing dimension $d$ and Weyl constant $\gamma_{d}$ to be recovered from one-dimensional bulk data. This transfer is stable under perturbations $\delta(\lambda)=o(\lambda)$ (Proposition 5.3), with error rates governed by the slowly varying factor $\eta=\delta/\lambda\in\mathrm{RV}_{0}$.
To make the rigidity statement structural (rather than only model-specific), we introduce asymptotic spectral equivalence classes determined by regular variation exponent and define the encoding-induced map on these classes. For $\phi\in\mathrm{RV}_{k}$, this map scales asymptotic spectral dimension by $d_{\mathrm{as}}\mapsto d_{\mathrm{as}}/k$, so the dimension-preserving case is exactly $\phi\in\mathrm{RV}_{1}$ (with strict affine normalization when $L(\lambda)\to 1$).
Beyond uniqueness and stability, we prove that the full discrete spectral measure with multiplicities — capturing clustering on symmetric spaces — is completely and uniquely realizable by a Kreĭn string (Theorem 6.5, [14]). No classical smooth Sturm–Liouville operator can do this for $d>1$ (Lemma 6.11, [17, 13]), so the Kreĭn model is the natural one-dimensional realization of Riemannian spectra.
Theorem 1.1 (Affine Rigidity Theorem).
Let $(M^{d},g)$ be a smooth compact Riemannian manifold without boundary.
  1. (i)
    (O-regular variation rigidity.) Among all O-regularly varying encodings $C=\pi-\phi(\lambda)$, the requirement that the bulk density exponent equals (d2)/2(d-2)/2 forces $\phi\in\mathrm{RV}_{1}$, i.e. asymptotic linearity of the encoding.
  2. (ii)
    (Polynomial special case.) In particular, among all polynomial-type encodings $C=\pi-\epsilon\lambda^{k}L(\lambda)$ with $k>0$ and $L\in\mathrm{RV}_{0}$, only the asymptotically affine case $k=1$ preserves the geometric bulk density exponent (d2)/2(d-2)/2. Any $k\neq 1$ distorts the exponent to $(d-2)/(2k)\neq(d-2)/2$, making spectral-dimension recovery impossible.
Corollary 1.2 (Edge-variable Weyl law and stability).
Under the affine encoding $C=\pi-\epsilon\lambda$:
  1. (i)
    (Weyl transfer.) In the bulk regime $C\to$-$\infty$,
@@
    enabling exact recovery of $d$ and $\gamma_{d}$ from one-dimensional encoded data in the bulk regime ($C\to$-$\infty$).
  2. (ii)
    (Stability.) The bulk exponents are preserved under perturbations $C=\pi-\epsilon\lambda+\delta(\lambda)$ with $\delta(\lambda)=o(\lambda)$, with error rates governed by the slowly varying factor $\eta=\delta/\lambda\in\mathrm{RV}_{0}$.
@@
The edge-variable framework was introduced informally in [1] (sec. VI.A), where spectral density in the edge variable was proposed as a dimensional signature. The present paper establishes the rigidity of this choice: the affine encoding is not merely convenient but is the unique polynomial-type reparametrization compatible with spectral-dimension recovery. The proofs rely on Tauberian tools from regular variation theory [9] and on the Kreĭn string correspondence [14].
@@
We reserve edge for the geometric endpoint $C\uparrow\pi$, while bulk refers to the high-energy regime $C\to$-$\infty$. Note that $C=\pi$ is the image of $\lambda_{0}=0$ under the encoding $C=\pi-\epsilon\lambda$: it is the spectral edge of the Riemannian Laplacian in the $C$-variable (the encoding reverses orientation, so the lower spectral boundary $\lambda=0$ maps to the upper boundary $C=\pi$). The terminology is inherited from the companion framework [1], where $C=\pi$ is the natural upper limit of the deformation operator $\hat{C}$ and the variable $C$ is called the edge variable; the Weyl-type asymptotics occur in the bulk ($C\to$-$\infty$) regime of this variable. For modern overviews of Weyl asymptotics and spectral Tauberian methods, see [22].
@@ 2 Preliminaries and standing assumptions
This section fixes notation and records the standing assumptions used throughout. We introduce the Laplace spectrum on a compact Riemannian manifold, the affine encoding $C=\pi-\epsilon\lambda$ and the induced edge-variable measures, the composition identity that connects edge-variable counting to the classical counting function, and the regular variation toolkit from which the uniqueness and stability arguments draw.
@@ 2.1 Geometry and Laplace spectrum
We work throughout on a smooth compact Riemannian manifold without boundary; the Laplace–Beltrami operator has discrete spectrum with finite multiplicities.
@@ 2.2 Spectral measures and counting
We encode the spectrum as a discrete measure and record the càdlàg properties of the associated counting function used throughout.
@@
$N_{\Delta}(\Lambda):=\mu_{\Delta}((-\infty,\Lambda])=\sum_{\lambda_{n}\leq\Lambda}m_{n},$ (2.3)
where eigenvalues are repeated according to their multiplicities $m_{n}$ [4, Thm. VII.2].
@@
Since $N_{\Delta}(\Lambda)=\mu_{\Delta}((-\infty,\Lambda])$ with a purely atomic measure, all stated properties are immediate. ∎
@@ 2.3 Affine edge encoding and edge variable
Immediate from $dC/d\lambda=-\epsilon<0$. ∎
@@ 2.4 Exact composition identity
The following lemma is the algebraic core of the paper: it shows that the edge-variable counting function is identical to the classical counting function evaluated at a rescaled argument, making every asymptotic result about $N_{\Delta}$ immediately available for $N_{\mu_{C}}$.
@@
Since $C=\pi-\epsilon\lambda$ is strictly decreasing, $\{n:C_{n}\geq C\}=\{n:\lambda_{n}\leq(\pi-C)/\epsilon\}$, with multiplicities. ∎
@@ 2.6 Standing assumptions
For the reader’s convenience we collect all hypotheses used in the paper into a single assumption block.
@@
with the classical remainder $O(\Lambda^{(d-1)/2})$ when available; [6, 20, 21, 7, 8] and modern expositions [22, 23, 25];
@@
3 Weyl Law in the Edge Variable
The rigidity theorem (Theorem 1.1) singles out the affine encoding as the unique preserver of the geometric density exponent. This section records the positive side of that rigidity: what the affine encoding actually delivers. The key identity $N_{\mu_{C}}(C)=N_{\Delta}((\pi-C)/\epsilon)$ (Lemma 2.7) reduces every asymptotic question about $N_{\mu_{C}}$ to the corresponding question about $N_{\Delta}$, so the classical Weyl law transfers immediately; the remainder estimate, uniformity in $\epsilon$, and the two-term boundary case are then recorded as separate propositions.
@@ 3.1 Edge-variable Weyl (Abelian direction)
The following theorem establishes the Weyl law in the edge variable and its converse; the converse is the reconstruction step that lets one read off $d$ and $\gamma_{d}$ from bulk edge-variable data.
@@
$F$ is monotone as a counting function; we apply the smoothing of Prop. 3.12 only to interpret $F^{\prime}$ as a classical derivative, without affecting the asymptotics. By Karamata’s differentiation (monotone density) theorem for regularly varying functions (Bingham–Goldie–Teugels [9, §§1.4–1.7, Thm. 1.7.2]; see also [22, 23]), if $F(y)=y^{\alpha}\ell(y)$ with $\alpha>0$ and $\ell\in\mathrm{RV}_{0}$ is ultimately monotone, then $F^{\prime}(y)\sim\alpha y^{\alpha-1}\ell(y)$. Hence
@@
so necessarily $\alpha=d/2$ and $\gamma_{d}=A\,\epsilon^{d/2}$. ∎
@@
with ultimate monotonicity (or after smoothing as in Prop. 3.12), then
@@
The log–log slope $\alpha=\lim_{y\to\infty}\frac{d\log N_{\mu_{C}}(\pi-y)}{d\log y}$ yields $d=2\alpha$, and
@@
The same holds using the bulk density $\rho_{\mathrm{bulk}}(C)\sim B\,y^{\alpha-1}$, where $d=2\alpha$ and $B=\tfrac{d}{2}\gamma_{d}\,\epsilon^{-d/2}$. In practice, given samples of $N_{\mu_{C}}(C)$ for large negative $C$, consistent estimators are
$\widehat{d}(C):=2\,\frac{d\log N_{\mu_{C}}(C)}{d\log(\pi-C)},\qquad\widehat{\gamma}_{d}(C):=\epsilon^{\widehat{d}(C)/2}\,\frac{N_{\mu_{C}}(C)}{(\pi-C)^{\widehat{d}(C)/2}};$ (3.15)
in discrete data, replace the derivative by a local log–log slope. Clustering contributes only bounded microscopic oscillations and does not affect the leading bulk asymptotics (see Lemma 6.14).
@@ 3.2 Weyl remainder in the edge variable (bulk regime)
The next propositions quantify how the Weyl remainder and the uniformity in $\epsilon$ transfer to the edge variable, and extend the main asymptotic to the smoothed density.
@@
If $N_{\Delta}(\Lambda)=\gamma_{d}\Lambda^{d/2}+O(\Lambda^{(d-1)/2})$ as $\Lambda\to\infty$ (see [6, 20, 21, 7, 8]), then, as $C\to$-$\infty$ (equivalently, $\Lambda=(\pi-C)/\epsilon\to\infty$),
@@
Apply Lemma 2.7 with $\Lambda=(\pi-C)/\epsilon$ and propagate the remainder. The substitution $\Lambda=(\pi-C)/\epsilon$ and monotonicity preserve both the order and form of the remainder. ∎
@@
since $N_{\Delta}(Y)\sim\gamma_{d}Y^{d/2}$. ∎
@@
If $N_{\Delta}(\Lambda)=\gamma_{d}\Lambda^{d/2}+O(\Lambda^{(d-1)/2})$ as $\Lambda\to\infty$, then after standard smoothing (as in Prop. 3.12), the bulk density satisfies
@@
Apply the monotone density theorem to the remainder estimate from Proposition 3.4, using the same smoothing procedure as in Proposition 3.12. ∎